A department store marks all of their men’s dress shirts at 15% off. What’s the new price in dollars and cents for a shirt that originally costs $25? Do not include the dollar sign in your answer, just the number in dollars and cents
Answer #1

Original Price of shirt = $25.

Store giving 15% off.

That is, store giving 25*0.15=$3.75 off.

So, new price of shirt = 25-3.75 = 21.25